OpenStudy (anonymous):

An airplane if flying at an airspeed of 510 km/hr in a wind blowing at 60 km/hr toward the southeast. In what direction should the plan head to end up going due east?

OpenStudy (anonymous):

@satellite73

OpenStudy (ddcamp):

 Do you know how to find the ?°

OpenStudy (anonymous):

Um, it should be arctan of 60/510?

OpenStudy (anonymous):

@DDCamp

OpenStudy (anonymous):

use sine rule

OpenStudy (anonymous):

How would that look?

OpenStudy (anonymous):

you can't use SOHCAHTOA because it's not a right triangle \[\frac{sinA}{a} = \frac{sinB}{b} = \frac{sinC}{c}\]